OpenStudy (specek18):

PLEASE HELP!!!!!!! WILL GIVE MEDALS:) complete this statement: 40x^4a^2+16x^2a^3=8x^2a^2( )

OpenStudy (specek18):

@AnswerMyQuestions @Albany_Goon @amandanewsomm @butterflydreamer @bazel_e @bohotness @ChiefArnav @Abhisar

OpenStudy (bohotness):

what do you think completes the statment?

OpenStudy (specek18):

4x^2 is a common factor....that's really all i have so far...

OpenStudy (anonymous):

40x^4 a^2 + 16x^2 a^3 = 8x^2 a^2 ( ) so now equate this to zero imean do this 40x^4 a^2 + 16x^2 a^3 - 8x^2 a^2 = 0

OpenStudy (anonymous):

now solve this equation

OpenStudy (specek18):

:)working on it....one moment:)

OpenStudy (specek18):

i have a question...why did you subtract the 8x^2?

OpenStudy (anonymous):

as its positive on the side it belongs so on changing the side it becames negative

OpenStudy (specek18):

okay...

OpenStudy (anonymous):

hint: Common factor: 8x^2a^2

OpenStudy (specek18):

not sure if this is right but here's where i'm at 10x^2a^24xa^3-4xa^2=0

OpenStudy (anonymous):

thats wrongs

OpenStudy (anonymous):

look at the hint i gave you and the equation

OpenStudy (specek18):

you can't do 8x because it's not a multiple of 40...

OpenStudy (specek18):

just kidding. i'm wrong. sorry!

OpenStudy (anonymous):

so what did you get now?

OpenStudy (anonymous):

8 times 5 is 40 -_-

OpenStudy (specek18):

10x^2a^24xa^3-4xa^2=0

OpenStudy (specek18):

and i know, horrible at math, sorry.

OpenStudy (anonymous):

scroll back up and look at the very first reply of mine...

OpenStudy (specek18):

i'm so confused now..i'm sorry just looked at it. what'd i do wrong?

OpenStudy (anonymous):

ummmm....? sorry to say but we will have begin all over again.. -_-

OpenStudy (specek18):

why?

OpenStudy (anonymous):

40x^4 a^2 + 16x^2 a^3 = 8x^2 a^2 ( ) so now equate this to zero imean do this 40x^4 a^2 + 16x^2 a^3 - 8x^2 a^2 = 0

OpenStudy (anonymous):

now take 8x^2 a^2 as a common factor

OpenStudy (specek18):

can you just give me the answer? i don't have enough time to do it again...it's a timed thing.

OpenStudy (specek18):

@bohotness do you know the answer?

OpenStudy (anonymous):

8x^2a^2[5x^2 + 2a - 1] is the answr

OpenStudy (anonymous):

answer*

OpenStudy (bohotness):

he is right:)

OpenStudy (bohotness):

nd yes i do known the answer

OpenStudy (specek18):

okay so thats wrong...the correct answer was 5x^2+2a but thank you for your help anyways.

OpenStudy (bohotness):

\[40x^{2}a ^{2}+16x^{2}a ^{3} divide 8x ^{2}a ^{2}